OpenStudy (anonymous):

If cosine 70 degrees= 0.342 give another angle whose cosine is 0.342??

OpenStudy (lgbasallote):

the coterminal angle...just add 360 to 70

OpenStudy (lgbasallote):

i think -70 also qualifies

OpenStudy (lgbasallote):

yupp cos(-70) qualifies :D

OpenStudy (anonymous):

What about cos x= -cos(180-x)

OpenStudy (campbell_st):

I think think you can use cos(x) = a then sin(90 - x) = a

OpenStudy (anonymous):

my does my text give the answer 260??

OpenStudy (campbell_st):

well cos is positive in the 1st and 4th quadrant... 4th quadrant angles are \[360 - \theta\] so 360 - 70 = 290 might me a textbook typo

OpenStudy (anonymous):

ya its 290